OpenStudy (anonymous):

collect like terms 16r+3s-6r-9s=

OpenStudy (anonymous):

16r-6r = 10r then 3s-9s = -6s so that will be: 10r-6s. Just add and subtract the numbers with the same letter. (aka simplify like terms) ^^

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

no. you don't subtract 10r from -6s. they have different variables (variables - the letter beside the number). so the final answer will be 10r-6s=0.

OpenStudy (anonymous):

Why does it have to equal 0?

OpenStudy (anonymous):

i didnt get that answer, i got 2

OpenStudy (anonymous):

to make it an equation form. but it can stand alone as just 10r-6s

OpenStudy (anonymous):

How did u get 2?

OpenStudy (anonymous):

please post your solution. i'll try to figure it out xD

OpenStudy (anonymous):

2(5r-3s)

OpenStudy (anonymous):

Simplifying 16r + 3s + -6r + -9s = 0 Reorder the terms: 16r + -6r + 3s + -9s = 0 Combine like terms: 16r + -6r = 10r 10r + 3s + -9s = 0 Combine like terms: 3s + -9s = -6s 10r + -6s = 0 Solving 10r + -6s = 0 Solving for variable 'r'. Move all terms containing r to the left, all other terms to the right. Add '6s' to each side of the equation. 10r + -6s + 6s = 0 + 6s Combine like terms: -6s + 6s = 0 10r + 0 = 0 + 6s 10r = 0 + 6s Remove the zero: 10r = 6s Divide each side by '10'. r = 0.6s Simplifying r = 0.6s

OpenStudy (anonymous):

yeah, your solution is correct ^^ and i was shocked. you didn't include finding the value of r to your question, lol XD